OpenStudy (anonymous):

Please help ASAP!! Ten 6-sided dice are rolled. What is the probability that exactly three of the dice show a 1?

OpenStudy (anonymous):

PLEASE HELP ASAP

OpenStudy (xapproachesinfinity):

hmm i'm not sure about this my guess is to use combination 10 choose 3 multiply that by the probability of getting a 1 per one die

OpenStudy (xapproachesinfinity):

\(\Large \binom{10}{3}*(1/6)\) this how you write it

OpenStudy (xapproachesinfinity):

@SithsAndGiggles

OpenStudy (anonymous):

@xapproachesinfinity that's almost right... We're dealing with two possible outcomes here - the outcome that we roll a particular number (in this case a 1) and the outcome that we DO NOT roll that number (any of 2,3,4,5,6). Hence we have a binomial distribution with \(p=\dfrac{1}{6}\) and \(n=10\). The desired probability is given by \[\binom{10}3\left(\frac{1}{6}\right)^{10−3}\left(\frac{5}{6}\right)^3\]

OpenStudy (xapproachesinfinity):

i see, you are right i didn't consider the outcome for not rolling a 1 now i understand it what i was missing @SithsAndGiggles thanks